# publications.bib

### v3

@@ 3. Mantovan’s Formula and the Hodge-Tate Period Morphism
as desired. ∎
@@ 4.1. The Spectral Action
Let $G/\mathbb{Q}_{p}$ be a quasi-split connected reductive group with a choice of Borel $B$ and maximal torus $T$ as before. We will work with $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$, the derived category of étale $\overline{\mathbb{F}}_{\ell}$-sheaves on the moduli stack of $G$-bundles. Our goal in this section will be to describe a localization $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}\subset\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ for $\mathfrak{m}\subset H_{K_{p}^{\mathrm{hs}}}$ a generic maximal ideal with associated semi-simple L-parameter $\phi_{\mathfrak{m}}$ in the case that $G$ is unramified. We will do this in slightly more generality using the spectral action [FS, Section X.2]. We assume that $\ell$ is very good as in [FS, Page 33], and consider the moduli stack $\mathfrak{X}_{\hat{G}}/\mathop{\rm Spec}{\overline{\mathbb{F}}_{\ell}}$ of $\overline{\mathbb{F}}_{\ell}$-valued Langlands parameters, as defined in [DH, Zhu1]. We let $\mathrm{Perf}(\mathfrak{X}_{\hat{G}})$ denote the derived category of perfect complexes on this stack, and we write $\mathrm{Perf}(\mathfrak{X}_{\hat{G}})^{BW_{\mathbb{Q}_{p}}^{I}}$ for the derived category of objects with a continuous $W_{\mathbb{Q}_{p}}^{I}$ action for a finite index set $I$, and $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})^{\omega}$ for the triangulated sub-category of compact objects in $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$. By [FS, Corollary X.I.3], there exists a $\overline{\mathbb{F}}_{\ell}$-linear action
@@
which, extending by colimits, gives
$\mathrm{Ind}(\mathrm{Perf}(\mathfrak{X}_{\hat{G}}))\rightarrow\mathrm{End}(\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell}))$
We recall the following basic properties of this action.
@@
    For $V=\boxtimes_{i\in I}V_{i}\in\mathrm{Rep}_{\overline{\mathbb{F}}_{\ell}}(\phantom{}^{L}G^{I})$, there is an attached vector bundle $C_{V}\in\mathrm{Perf}(\mathfrak{X}_{\hat{G}})^{BW_{\mathbb{Q}_{p}}^{I}}$ whose evaluation at a $\overline{\mathbb{F}}_{\ell}$-point of $\mathfrak{X}_{\hat{G}}$ corresponding to a (not necessarily semi-simple) $L$-parameter $\tilde{\phi}:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}G(\overline{\mathbb{F}}_{\ell})$ is the vector space $V$ with $W_{\mathbb{Q}_{p}}^{I}$-action given by $\boxtimes_{i\in I}r_{V_{i}}\circ\tilde{\phi}$. The endomorphism
@@
    The action is monoidal in the sense that, given $C_{1},C_{2}\in\mathrm{Perf}(\mathfrak{X}_{\hat{G}})$, we have a natural equivalence of endofunctors
@@
If $\phi:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}G(\overline{\mathbb{F}}_{\ell})$ is a semi-simple $L$-parameter then this defines a closed $\overline{\mathbb{F}}_{\ell}$-point $x$ in the moduli stack of Langlands parameters, which maps to a closed point in the coarse moduli space. We let $\mathfrak{m}_{\phi}\subset\mathcal{O}_{\mathfrak{X}_{\hat{G}}}(\mathfrak{X}_{\hat{G}})$ denote the corresponding maximal ideal. We recall that, for all $f\in\mathcal{O}_{\mathfrak{X}_{\hat{G}}}(\mathfrak{X}_{\hat{G}})$ and $A\in\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$, one obtains an endomorphism
@@
induced by multiplication by $f$. Under the description of $\mathcal{O}_{\mathfrak{X}_{\hat{G}}}(\mathfrak{X}_{\hat{G}})$ in terms of the excursion algebra, this encodes the action of the excursion algebra on $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ [Zou, Theorem 5.2.1]. More precisely, we recall that, since $\ell$ is very good [FS, Page 33], to any Schur-irreducible $A\in\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ we can, by [FS, Proposition I.9.3], attach a conjugacy class of semi-simple $L$-parameters
@@
called the Fargues-Scholze parameter of $A$. By [FS, Theorem VIII.3.6], we have an identification between the ring of global functions $\mathcal{O}_{\mathfrak{X}_{\hat{G}}}(\mathfrak{X}_{\hat{G}})$ and excursion operators. Since $A$ is Schur irreducible the endomorphisms corresponding to $f\in\mathcal{O}_{\mathfrak{X}_{\hat{G}}}(\mathfrak{X}_{\hat{G}})$ determine a non-zero scalar in $\overline{\mathbb{F}}_{\ell}$ which will be determined by the excursion datum evaluated at the Fargues-Scholze parameter $\phi_{A}^{\mathrm{FS}}$.
@@
We define $\iota_{\phi}:\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}\hookrightarrow\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ to be the full-subcategory of objects $A$ for which the endomorphisms $A\rightarrow A$ induced by $f\in\mathcal{O}_{\mathfrak{X}_{\hat{G}}}\setminus\mathfrak{m}_{\phi}$ are isomorphisms.
It is easy to check that the subcategory $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}\subset\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ is preserved under colimits and limits, and therefore, by the $\infty$-categorical adjoint functor theorem [HTT, Corollary 5.5.2.9], there exists a left adjoint to the inclusion $\iota_{\phi}$ denoted by $\mathcal{L}_{\phi}$. We define $(-)_{\phi}:=\iota_{\phi}\mathcal{L}_{\phi}(-)$. This, by the fully faithfulness of $\iota_{\phi}$, will define an idempotent functor (See Appendix A for details).
@@
    Given $A\in\mathrm{D}(G(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})\subset\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$, we have an isomorphism
@@
    where the LHS is the usual localization under the smooth Hecke algebra.
@@
    If $A\in\mathrm{D}_{\mathrm{lis}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ is ULA then one has a direct sum decomposition
@@
For (1), this follows since the action of $\mathcal{O}_{\mathfrak{X}_{\hat{G}}}(\mathfrak{X}_{\hat{G}})$ on $A$ will factor through the maximal ideal $\mathfrak{m}_{A}$ defined by the semi-simple $L$-parameter $\phi_{A}^{\mathrm{FS}}$ attached to $A$ by the above discussion, and therefore $A\in\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$ forces an equality of maximal ideals: $\mathfrak{m}_{A}=\mathfrak{m}_{\phi}$.
For (3), we use the arguments in Koshikawa [Ko, Page 6]. Consider the map
@@
given by the spectral action, where $\mathrm{cInd}_{K_{p}^{\mathrm{hs}}}^{G(\mathbb{Q}_{p})}(\overline{\mathbb{F}}_{\ell})$ is regarded as a right $H_{K_{p}^{\mathrm{hs}}}$-module. It is shown that after through the usual action by the unramified Hecke algebra composed with the involution $KhK\rightarrow Kh^{-1}K$ gives rise to a map which is compatible with usual $L$-parameters for unramified irreducible representations. In particular, the pullback of the maximal ideal $\mathfrak{m}\subset H_{K_{p}^{\mathrm{hs}}}$ is given by the maximal ideal $\mathfrak{m}_{\phi_{\mathfrak{m}}}\subset\mathcal{O}_{X_{\hat{G}}}(\mathfrak{X}_{\hat{G}})$. Now, by arguing as in Proposition A.3, we have an identification:
@@
Let $A$ be a complex of smooth $G(\mathbb{Q}_{p})$-representations which is admissible (i.e $A^{K}$ is a perfect complex for all compact open $K\subset G(\mathbb{Q}_{p})$). We then have a decomposition
@@
running over semisimple $L$-parameters, where any irreducible constituent $\pi$ of $A_{\phi}$ has Fargues-Scholze parameter equal to $\phi_{\pi}^{\mathrm{FS}}$, as conjugacy classes of parameters.
@@
This follows by applying to Lemma 4.2 (1) and (4) to the full subcategory $\mathrm{D}(G(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})\subset\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$
Now our goal is to describe the subcategory $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$ more explicitly, using the results of [Ham2] in the case that $\phi$ is induced from a generic toral parameter $\phi_{T}$. To do this, we will need to have some information about the Fargues-Scholze local Langlands correspondence. First, let us introduce some notation.
We let $B(G)_{\mathrm{un}}:=\text{Im}{(B(T)\rightarrow B(G))}$. We recall that these are precisely the elements $b\in B(G)$ such that the $\sigma$-centralizer $J_{b}$ is quasi-split ([Ham2, Lemma 2.12]). In particular, the fixed choice of Borel $B\subset G$ transfers to a Borel subgroup $B_{b}$ for all $b\in B(G)_{\mathrm{un}}$, and $J_{b}\simeq M_{b}$ under the inner twisting, where $M_{b}\subset G$ is the Levi subgroup of $G$ determined by the centralizer of the slope homomorphism of $b$ in $G$. We let $\delta_{P_{b}}$ denote the modulus character of the standard parabolic $P_{b}$ of $G$ with Levi factor $M_{b}$ transferred to $J_{b}$ under the inner twisting. We set $W_{b}:=W_{G}/W_{M_{b}}$ to be the quotient of the relative Weyl group of $G$ by the relative Weyl group of $M_{b}$. We identify $W_{b}$ with a choice of representatives in $w\in W_{G}$ of minimal length. We set $\rho_{b,w}:=i_{B_{b}}^{J_{b}}(\chi^{w})\otimes\delta_{P_{b}}^{-1/2}$ to be the normalized parabolic induction of $\chi^{w}$, where $\chi$ is the character of $T(\mathbb{Q}_{p})$ attached to a toral parameter $\phi_{T}$ under local class field theory and $\delta_{P_{b}}$ is the modulus character of $M_{b}\simeq J_{b}$.
We will need to assume the following properties of the Fargues-Scholze local Langlands correspondence, as in [Ham2, Assumption 7.5].

### v5

@@ 3.2. Mantovan’s Product Formula
as desired (See [25, Corollary 1.9 (1)]). ∎
Remark 3.23.
The analogous statement involving the full Hecke operator $T_{\mu}:\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})\rightarrow\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})^{BW_{E_{\mu}}}$ and the natural Weil group action on $R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C})$ also holds. The process of matching the actions of the inertia group $I_{E_{\mathfrak{p}}}\subset W_{E_{\mathfrak{p}}}$ is fairly straight forward; in particular, one just needs to check that all the above Cartesian diagrams descend to $\breve{E}_{\mathfrak{p}}$. The process of matching the Frobenius descent datum is a bit tricky however so we do not address this here; for a relevant discussion of this see [10, Sections 8.2,8.4.9].
@@ 4. The Local Results
In this section, we establish the key local results (Corollaries 4.27,4.29) necessary for the proof of Theorem 1.8. We will let $G/\mathbb{Q}_{p}$ be a quasi-split connected reductive group with a choice of Borel $B$ and maximal torus $T$ as before. For the application of these local results to Theorem 1.8, we will in practice take $G$ to be the base-change of the global group $\mathbf{G}$ to $\mathbb{Q}_{p}$.
More specifically, for a semi-simple $L$-parameter $\phi$, in §4.1, we review the spectral action of the category of quasi-coherent sheaves on the stack of L-parameters on $\mathrm{D}(\mathrm{Bun}_{G})$, as well as its relationship to the action of Hecke operators on $\mathrm{D}(\mathrm{Bun}_{G})$. This will allow us to introduce a certain localization functor $\mathrm{D}(\mathrm{Bun}_{G})\rightarrow\mathrm{D}(\mathrm{Bun}_{G})_{\phi}$ that Hansen will study more extensively in Appendix A. We establish some basic properties of the localization functor that will be used in the later sections, and then turn our attention to studying the derived category $\mathrm{D}(\mathrm{Bun}_{G})_{\phi}$ for $\phi$ induced from a generic toral L-parameter. In particular, in §4.2, we discuss how, assuming the Fargues-Scholze local Langlands correspondence is compatible with a suitably nice form of local Langlands, one can show that every object in $\mathrm{D}(\mathrm{Bun}_{G})_{\phi}$ can be described in terms of constituents of certain explicit parabolic inductions. We then collect in §4.2.2 various instances for which this compatibility holds, using the results of [47, 24, 31].
In §4.3, we turn to the perverse $t$-exactness of the Hecke operators on the subcategory $\mathrm{D}(\mathrm{Bun}_{G})_{\phi}$, explaining how the theory of geometric Eisenstein series allows one to verify the perverse t-exactness for the subcategory spanned by the inductions described in §4.2, under some possible constraints including the genericity and regularity of a toral parameter $\phi_{T}$ inducing $\phi$. This will allow us to prove our main results under some possible additional constraints. In §4.4, we study the relationship between the genericity and these additional constraints, for the groups for which compatibility of the Fargues-Scholze local Langlands is known, showing that the additional constraints always follow from genericity in the relevant cases. This will allow us to conclude the precise form of our main local results.
@@ 4.1. The Spectral Action
[…]
@@
which, extending by filtered colimits, gives
$\mathrm{Ind}(\mathrm{Perf}(\mathfrak{X}_{\hat{G}}))\rightarrow\mathrm{End}(\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})),$
where the notation $\mathrm{Ind}(-)$ is the operation described in [45, Section 5.3]. We recall the following basic properties of this action [^9] These properties follow from the construction, since it is given by [16, Corollary X.I.3] applied to the functorial in $I$ Hecke action of $\mathrm{Rep}_{\overline{\mathbb{F}}_{\ell}}(\phantom{}^{L}G^{I})$ on $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})^{\omega}$, as at the end of [16, Section X.3]. In particular, property (1) immediately follows from this, and property (2) also follows since the action appearing in [16, Corollary X.I.3] is monoidal..
@@
    For $V=\boxtimes_{i\in I}V_{i}\in\mathrm{Rep}_{\overline{\mathbb{F}}_{\ell}}(\phantom{}^{L}G^{I})$, there is an attached vector bundle $C_{V}\in\mathrm{Perf}(\mathfrak{X}_{\hat{G}})^{BW_{\mathbb{Q}_{p}}^{I}}$ whose pullback to a $\overline{\mathbb{F}}_{\ell}$-point of $\mathfrak{X}_{\hat{G}}$ corresponding to a (not necessarily semi-simple) $L$-parameter $\tilde{\phi}:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}G(\overline{\mathbb{F}}_{\ell})$ is the vector space $V$ with $W_{\mathbb{Q}_{p}}^{I}$-action given by $\boxtimes_{i\in I}r_{V_{i}}\circ\tilde{\phi}$, where $r_{V_{i}}:\phantom{}^{L}G\rightarrow\mathrm{GL}(V_{i})$ is morphism defined by $V_{i}$. More specifically, these bundles are given by pulling back $V|_{\hat{G}^{I}}$ along the natural map $\mathfrak{X}_{\hat{G}}\rightarrow[\mathop{\rm Spec}(\overline{\mathbb{F}}_{\ell})/\hat{G}$], where we note that these have a natural $W_{\mathbb{Q}_{p}}^{I}$-equivariant structure coming from the full action of $\phantom{}{}^{L}G$ on $V$. The endomorphism
@@
[…]
@@
Remark 4.1.
The Hecke operators appearing here are slightly different then the ones discussed in §3.2. In particular, for $\mu$ a geometric dominant cocharacter of $G$, we recall that there is a canonical away of attaching a representation of $\phantom{}{}^{L}G$. Namely, by the procedure of [38, Lemma 2.1.2], the highest weight representation $V_{\mu}$ of $\hat{G}$ descends to a representation of $\hat{G}\rtimes W_{E_{\mu}}$ where $E_{\mu}$ is the reflex field of the cocharacter $\mu$. This can then be induced along the inclusion $W_{E_{\mu}}\subset W_{\mathbb{Q}_{p}}$ to get a full representation of $\phantom{}{}^{L}G$ which we denote by $V$. Correspondingly, there is a Hecke operator
$T_{V}:\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})\rightarrow\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})^{BW_{\mathbb{Q}_{p}}}$
attached to the representation of $\phantom{}{}^{L}G$. This Hecke operator factors as
composed with the map $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})^{BW_{E_{\mu}}}\rightarrow\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})^{BW_{\mathbb{Q}_{p}}}$ induced by induction along the inclusion $W_{E_{\mu}}\subset W_{\mathbb{Q}_{p}}$ (cf. the discussion on [16, Page 322]). The Hecke operator $T_{\mu}:\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})\rightarrow\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})^{BW_{E_{\mu}}}$ was the one appearing in §3.2. We will often cite results for Hecke operators $T_{V}$ for $V\in\mathrm{Rep}_{\overline{\mathbb{F}}_{\ell}}(\phantom{}^{L}G)$ to deduce results about the $T_{\mu}$, implicitly using this dictionary in the process.
[…]
@@
[…]
@@
[…]
@@
We define $\iota_{\phi}:\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}\hookrightarrow\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ to be the full-subcategory of objects $A$ for which the endomorphisms $A\rightarrow A$ induced by $f\in\mathcal{O}_{\mathfrak{X}_{\hat{G}}}(\mathfrak{X}_{\mathfrak{X}_{\hat{G}}})\setminus\mathfrak{m}_{\phi}$ are isomorphisms.
It is easy to check that the subcategory $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}\subset\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ is preserved under colimits and limits, and therefore, by the $\infty$-categorical adjoint functor Theorem [45, Corollary 5.5.2.9], there exists a left adjoint to the inclusion $\iota_{\phi}$ denoted by $\mathcal{L}_{\phi}$ (See Appendix A for details). We define $(-)_{\phi}:=\iota_{\phi}\mathcal{L}_{\phi}(-)$. This, by the fully faithfulness of $\iota_{\phi}$, will define an idempotent functor.
@@
    Given $A\in\mathrm{D}(G(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$, we have an isomorphism
@@
    where the LHS is the usual localization under the smooth unramified Hecke algebra $H_{K_{p}^{\mathrm{hs}}}$ at some maximal ideal $\mathfrak{m}\subset H_{K_{p}^{\mathrm{hs}}}$. Here we abuse notation and identify $A\in\mathrm{D}(G(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$ with its image under the fully faithful embedding $\mathrm{D}(G(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})\xrightarrow{j_{1!}}\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$.
@@
    If $A\in\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ is ULA then one has a direct sum decomposition
@@
[…]
For (3), we use the arguments in Koshikawa [37, Page 6]. Consider the map
@@
[…]
@@
Let $A$ be a complex of smooth $G(\mathbb{Q}_{p})$-representations which is admissible (i.e $A^{K}$ is a perfect complex for all open pro-$p$ $K\subset G(\mathbb{Q}_{p})$). We then have a decomposition
@@
running over semisimple $L$-parameters, where any irreducible constituent $\pi$ of $A_{\phi}$ has Fargues-Scholze parameter equal to $\phi$, as conjugacy classes of parameters.
@@
This follows by applying to Lemma 4.3 (1) and (4) to the full subcategory $\mathrm{D}(G(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})\subset\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$, where we recall that a sheaf $A\in\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ lies in the ULA subcategory if and only if its pullbacks $j_{b}^{*}(A)\in\mathrm{D}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})\simeq\mathrm{D}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$ for all $b\in B(G)$ are admissible as in the statement of the Corollary ([16, Theorem V.7.1]) . ∎
We now turn our attention to studying the subcategory $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ in certain cases coming from generic toral parameters, by using compatibility of the Fargues-Scholze local Langlands correspondence with more classical and well understood instances of local Langlands.
4.2. Local-Global Compatibility of the Fargues-Scholze Local Langlands
Now our goal is to describe the subcategory $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$ more explicitly, using the results of [24] in the case that $\phi$ is induced from a generic toral parameter $\phi_{T}$, as in Definition 1.1. To do this, we will need to have some information about the Fargues-Scholze local Langlands correspondence. We first introduce the main result in this direction. We keep our running assumption that $\ell\nmid|\pi_{0}(Z(G))|$, so that the localized category $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$ described in the previous section is well-defined.
4.2.1. Implications of Compatibility for Generic Parameters
We let $B(G)_{\mathrm{un}}:=\text{Im}{(B(T)\rightarrow B(G))}$. We recall that these are precisely the elements $b\in B(G)$ such that the $\sigma$-centralizer $J_{b}$ is quasi-split ([24, Lemma 2.12]). In particular, the fixed choice of Borel $B\subset G$ transfers to a Borel subgroup $B_{b}$ for all $b\in B(G)_{\mathrm{un}}$, and $J_{b}\simeq M_{b}$ under the inner twisting, where $M_{b}\subset G$ is the Levi subgroup of $G$ determined by the centralizer of the slope homomorphism of $b$ in $G$. We let $\delta_{P_{b}}$ denote the modulus character of the standard parabolic $P_{b}$ of $G$ with Levi factor $M_{b}$ transferred to $J_{b}$ under the inner twisting (See [25, Definition 3.14] for a more precise discussion, where we use the same conventions for the modulus character as loc.cit). We set $W_{b}:=W_{G}/W_{M_{b}}$ to be the quotient of the relative Weyl group of $G$ by the relative Weyl group of $M_{b}$. We identify $W_{b}$ with a choice of representatives in $w\in W_{G}$ of minimal length. We set $\rho_{b,w}:=i_{B_{b}}^{J_{b}}(\chi^{w})\otimes\delta_{P_{b}}^{-1/2}$ to be the normalized parabolic induction of $\chi^{w}$, where $\chi$ is the character of $T(\mathbb{Q}_{p})$ attached to a toral parameter $\phi_{T}$ under local class field theory and $\delta_{P_{b}}$ is the modulus character of $M_{b}\simeq J_{b}$ (Here we have slightly abused notation by omitting $\chi$ from the notation, where it will usually be clear from context. If further clarifications is necessary then we will write $\rho_{b,w}^{\chi}$.).
We will need to assume the following properties of the Fargues-Scholze local Langlands correspondence, as in [24, Assumption 6.5].